OpenStudy (anonymous):

help y-(-2)= -3/4(x-2)

TheSmartOne (thesmartone):

\(\Huge\color{Blue}{Welcome~to~Open~Study!!}\)

TheSmartOne (thesmartone):

What are you solving for...

TheSmartOne (thesmartone):

@kdblankenship48

OpenStudy (anonymous):

Putting it in standard form ax+by=c

TheSmartOne (thesmartone):

So first we need to take care of y-(-2) then we need to distribute the -3/4 into the parenthesis.

TheSmartOne (thesmartone):

\[a-(-b)=a+b\] so what does y-(-2)=??

OpenStudy (anonymous):

lol can you explain it a little simpler, im a junior in highschool.

TheSmartOne (thesmartone):

well so am I...

OpenStudy (anonymous):

well, im struggling and you seem a lot smarter.

TheSmartOne (thesmartone):

When you subtract a negative number you have to add them together...

OpenStudy (anonymous):

i got that part

TheSmartOne (thesmartone):

so a-(-b)=a+b and y-(-2)=??? a=y b=2 Then y+2 So y-(-2)=y+2

OpenStudy (anonymous):

lol i was gonna put that but i thought i was wrong

TheSmartOne (thesmartone):

Always believe in yourself. If you get it wrong.. I am here to help you understand...

OpenStudy (anonymous):

so y+2=-3/4x-6/8?

TheSmartOne (thesmartone):

umm.. no

OpenStudy (anonymous):

ughhhh

TheSmartOne (thesmartone):

Ok so we got the y+2 lets leave that alone. Lets distribute the -3/4 inside...

OpenStudy (anonymous):

okay

TheSmartOne (thesmartone):

\[\frac{ -3 }{ 4 }(x-2)=\frac{ -3 }{ 4 }x -(2)(\frac{ -3 }{4 })\]

TheSmartOne (thesmartone):

Distributing inside means multiplying it inside to each value...

TheSmartOne (thesmartone):

\[a(x+y)=ax+ay\]

OpenStudy (anonymous):

i need to get the equation to y=mx+b and then change it to standard form

TheSmartOne (thesmartone):

So you should get so far \[y+2=\frac{ -3 }{ 4 }x+\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

yaaaaassss thats what i got

TheSmartOne (thesmartone):

Because \[\frac{ 3 }{ 4 }\times 2=\frac{ 3 }{ 2 }\times 1\]

TheSmartOne (thesmartone):

\[y+2-2=\frac{ -3 }{ 4 }x+\frac{ 3 }{ 2 }-2\]

TheSmartOne (thesmartone):

So what does y=??

OpenStudy (anonymous):

okay y=-3/4x-1/2?

TheSmartOne (thesmartone):

\[y=-\frac{ 3 }{ 4 }x-\frac{ 1 }{2 }\] is correct. ^_^

OpenStudy (anonymous):

yay! okay, now what?

TheSmartOne (thesmartone):

we just add 3/4x to get it into Ax+By=C

OpenStudy (anonymous):

so 3/4x+2y?

TheSmartOne (thesmartone):

No it would just be 3/4x+y=-1/2

OpenStudy (anonymous):

oh okay

TheSmartOne (thesmartone):

There is an implied 1 in front of the y because 1y=y, 1x=x, 1a=a

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

so that's the answer?

TheSmartOne (thesmartone):

Yup...

TheSmartOne (thesmartone):

You are new here right..

OpenStudy (anonymous):

lol yeah, my grandma was trying to help me so she went and made this account under her name and told me to get someone to help me, so here i am.

TheSmartOne (thesmartone):

Here is this tutorial I made for new people on OS. You can learn about more things you can do on OS. ^_^ http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

It doesn't really require a name.. Just an email address..

TheSmartOne (thesmartone):

yes..

TheSmartOne (thesmartone):

Wait we can have his talk in the messages...

OpenStudy (anonymous):

yeaaaah okay

TheSmartOne (thesmartone):

Do you know how to close this question?

TheSmartOne (thesmartone):

Here are some ways you can close it. http://prntscr.com/4wljgx http://prntscr.com/4wljjv

TheSmartOne (thesmartone):

@kdblankenship48